OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 7 inches and a base that is an equilateral triangle with sides of 6 inches?

OpenStudy (anonymous):

Volume will be the area of the eq triangle times the height.

OpenStudy (callisto):

First, find the area of the base By heron's formula area=sqrt 243 Volume = height x base area = sqrt243 x 7 = can you do it?

OpenStudy (callisto):

Heron's formula: \[Area \ of \ a \ triangle=\sqrt{s(s-a)(s-b)(s-c)}\], where s = (a+b+c)/2 a, b, c are the length of the sides of the triangle

OpenStudy (anonymous):

i still dont understand.

OpenStudy (callisto):

Which part?